Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2011-225_2011-231

Mission Operations Days: 2011/225 through 2011/231
August 13, 2011 20:00:00 GMT to August 19, 2011 20:00:00 GMT

Terra Spacecraft and MODIS Instrument Status:

AM-1 "Terra" is in Normal Mode
PFM MODIS is in A-Side Nominal Science Mode
(Cross Strapped with FR B)
ComponentState(s)Comment
BlackbodyA On (290K); B OffNominal Science Mode
Calibration ElectronicsA On; B OffNominal Science Mode
Control ProcessorA On; B OffNominal Science Mode
Nadir DoorUnlatched, OpenNominal Science Mode
Space View DoorUnlatched, OpenNominal Science Mode
Solar Diffuser DoorUnlatched, ScreenedPermanent Configuration
FDDI FormatterA On; B OffNominal Science Mode
FDDI PortA On; B OffNominal Science Mode
FIFO Memory#1 & #2 On; #3 & #4 OffNominal Science Mode
Format ProcessorA Off; B OnNominal Science Mode
Power Supplies#1 On, #2 OffNominal Science Mode
PV VIS FPAA On; B OffNominal Science Mode
PV NIR FPAA On; B OffNominal Science Mode
PV SMIR FPAA On; B OffNominal Science Mode
PV LWIR FPAA On; B OffNominal Science Mode
PC LWIR FPAA On; B OffNominal Science Mode
RC Outgas HeatersAll 3 (CS, IS, OS) OffNominal Science Mode
RC LWIR FPA HeaterOn (83K)Nominal Science Mode
RC SMIR FPA HeaterOffNominal Science Mode
Scan AssemblyA On; B OffNominal Science Mode
SDSMOffNominal Science Mode
SRCAOff
  • 10W Lamps - 1, 4 (functional)
  • 10W Lamps - 2, 3 (non-functional)
  • 1W Lamps - 1, 2 (functional)
Nominal Science Mode
PS1 Survival HeaterEnabledNominal Science Mode
PS2 Survival HeaterEnabledNominal Science Mode
Timing GeneratorA On; B OffNominal Science Mode
Flight SoftwareRev BD + 11 sets of patches
  • CP07 to Survival Mode - No SDD to closed command (1 patch)
  • CP07 to Safe Mode - No FR OFF Command (1 patch)
  • CP07 to Safe Mode - No SDD to closed command (1 patch)
  • CP07 to Science Mode - Turn on FRB instead of FRA (1 patch)
  • CP07 to Outgas and Science Mode - Turn on FRB instead of FRA (1 patch)
  • FR Error Counting and Statistics (2 patches)
  • FR 1000ms time code patches (3 patches)
  • FR Band 5 Gain Change (20 commands)
  • FR BB Ideal Output Values (1 patch)
  • FR A EEPROM KME Fast 1 word patch (1 patch)
  • FR B EEPROM KME Fast 1 word patch (1 patch)
Nominal Science Mode
Inhibit Ids SetnoneNominal Science Mode
TMON 66DisabledNominal Science Mode
TMON 67DisabledNominal Science Mode
TMON 68EnabledNominal Science Mode
TMON 69DisabledNominal Science Mode

This Week's Completed MODIS Activities:

Saturday, August 13, 2011

None

Sunday, August 14, 2011

None

Monday, August 15, 2011

None

Tuesday, August 16, 2011

None

Wednesday, August 17, 2011

None

Thursday, August 18, 2011

Lunar Roll Maneuver (-7.9336 degrees)
2011/230 16:06:19 ATC MOD_OA03_FR_NIGHT_RATE.1 (03:23 early)
2011/230 16:39:41-16:46:27 ATC MOD_OA08_ROLL_MAN.3
2011/230 17:02:24 ATC MOD_OA03_FR_DAY_RATE.2 (03:23 late)

Friday, August 19, 2011

None

This Week's Scheduled MODIS Activities Not Completed:

Sunday, August 14, 2011

2011/226 20:45:07-20:47:38 ATC MOD_OA27_PV_ECAL.3
OA-27: Photovoltaic bands electronics calibration.
Centered on S/C maximum latitude at 20:46:22.

Monday, August 15, 2011

2011/227 20:33:43-20:34:12 ATC MOD_SRCA_9_8_2_SETUP_CC.1
SRCA setup activity using 10W lamps 1 and 4, and 1W lamp 5, in constant
current feedback mode.
2011/227 20:36:43-21:08:08 ATC MOD_OA19_SRCA_FULL_RAD.1
OA-19: SRCA Full Radiometric Calibration.
Data collection 20:52:44-21:08:07 (centered on S/C Night)
S/C Night 20:42:59-21:16:53.

Tuesday, August 16, 2011

2011/228 20:21:05-20:29:53 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4
OA-16: SD/SDSM screened side A, no door movement.

This Month's Upcoming MODIS Events:

Calibrations for the end of July and August will include a Black Body Cycle (DOY 208-211), an SRCA Spectral (DOY 222), a PV ECAL (DOY 226), an SRCA Radiometric (DOY 227), an SDSM Screened (DOY 228), a Lunar Roll (DOY 230), a PV Ecal (DOY 233), and an SDSM Screened (DOY 235). Sunday, August 21, 2011

2011/233 20:51:15-20:53:46 ATC MOD_OA27_PV_ECAL.3
OA-27: Photovoltaic bands electronics calibration.
Centered on S/C maximum latitude at 20:52:30.

Tuesday, August 23, 2011

2011/235 20:27:53-20:36:44 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4
OA-16: SD/SDSM screened side A, no door movement.

Maneuvers:

None

MODIS Anomalies:

The MODIS IOT had to waive parts 2 - 4 of the scheduled SRCA Spectral activities on DOY 222. On two separate occasions during the scheduled contacts from 222/13:04-13:44 and 222/15:00-15:20, we attempted to load and dump the part 2 commands into macro 31 on board MODIS. Both times it seemed as if the files were loaded and dumped correctly, however during verification we noticed that on both occasions only 25 of the 26 files made it to the ground successfully. After double checking the logs and verifying that the scripts that were sent to the spacecraft were correct, we were unable to pin down the cause of the dump failures and were forced to waive the remainder of the activities due to time constraints. The failure to complete the remaining parts of the Spectral activities entailed us keeping inhibit IDs 33 and 54 set to prevent the activities from executing.

Due to inhibit ID 33 being set, all macros on boar MODIS will not execute. As a result, the scheduled PV Ecal activities on DOY 226 and the SDSM Screened activities on DOY 228 did not execute as scheduled. Inhibit ID 33 was cleared on DOY 230 prior to the Lunar Roll maneuver and the PV Ecal and SDSM activities were rescheduled (see above).

Also, as a result of the SRCA being left out of its home position after the failed execution of parts 2 - 4 of the Spectral activities on DOY 222, the scheduled SRCA Radiometric on DOY 227 had to be waived until the SRCA is brought back to its nominal home position which is scheduled to occur on DOY 238 (8/26/11).

General Instrument Comments:

MODIS is in Science Mode, using all A-side components, but cross-strapped to Formatter B, with the blackbody maintained at 290K and the Solar Diffuser Door at the Screened position.

MODIS Telemetry Trends:

Nominal telemetry trending is ongoing.

Non-MODIS Significant Events:

The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links web pages. There are directions at the URL below for how to obtain a username/password.

Following the Inclination Maneuver on DOY 286 (10/13/09), Terra experienced a Hex Bay Battery (BBAT) cell #50 failure and simultaneously (or close to same time) experienced a Battery Heater Control Electronic anomaly which prevents controlling the temperature for 5 of the 9 heater groups within BBAT. Cell #50 has failed and is non-recoverable. Battery temperatures have stabilized and battery cell temperatures are within an acceptable range. Terra has a robust power system with two 54 cell batteries that feed a power control system that maintains a stable 120V spacecraft power bus even if multiple cells fail. Terra is capable of continuing nominal operations with multiple cell failures. Science data acquisition has remained nominal throughout the anomaly and troubleshooting timeframe. Information will be updated periodically. At this point in time the Terra Battery Anomaly is not affecting science data collection and all MODIS systems are nominal. There has been little if any impact to operations since the issue began. All operations are nominal.

http://jupiter02.gsfc.nasa.gov/

Limited Life Item Status:

SRCA 10W Lamp #1: 298.8 of 500 hours
SRCA 10W Lamp #2: 172.1 of 500 hours
SRCA 10W Lamp #3: 190.3 of 500 hours
SRCA 10W Lamp #4: 100.2 of 500 hours

SRCA 1W Lamp #1: 581.7 of 4000 hours
SRCA 1W Lamp #2: 282.0 of 4000 hours

Solar Diffuser Door: 2146 of 3022 Movements
Nadir Aperture Door: 540 of 1316 Movements
Space View Door: 443 of 1316 Movements